Hello, I need help on a rational equation problem. 5/4 = n-3/n-4 I do realize that you either have to turn it into 4(5/4 = n-3/n-4) or n-4(5/4 = n-3/n-4) but I am sadly unsure of exactly how to do this correctly. I thought perhaps it would turn into 5 = 4n-3/4n-4 but that doesn't seem correct to me. Nor does (n-4)(5/4) = n-3 I'm really struggling with these and I have a test on Friday. Thank you for your help.
Answer by nerdybill(7384)

Think "cross-multiply"
5/4 = (n-3)/(n-4)
5(n-4) = 4(n-3)
5n-20 = 4n-12
n-20 = -12
n = 8